OpenStudy (anonymous):

Explain, in complete sentences and in your own words, the answers to the following questions relating to complex numbers. What are the characteristics of a radical equation? How is solving radical equations similar to solving linear equations? Why is it important to check the solutions to a radical equation? Create your own radical equation. Describe in complete sentences and demonstrate the process in finding its solution(s).

OpenStudy (kainui):

Sounds pretty radical, dude.

OpenStudy (anonymous):

A radical equation is when a variable is present in the form of a root. For example, the square-root of x + 8 = 12. You solve them like a linear equation because you still have to try to isolate the variable by removing all the other operations around it. Because often when you solve a radical equation, you end up taking the square along the way. This can introduce extraneous solutions. x + 2 = y Using the property of equality, I'm going to subtract 2 from both sides of the equal operator. x + 2 - 2 = y - 2 Now, I'm going to simplify by combining like terms. x = y - 2

OpenStudy (anonymous):

you are my savior <3

OpenStudy (anonymous):

♥ Xo!